# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

### v2

@@ Abstract
The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and underutilized spectral complementarity. Existing methods often fail to decouple shared structural features from modality-complementary radiometric attributes, causing feature conflicts and information loss. To address this, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-complementary) components in the Fourier domain, thus reinforcing shared structures while preserving complementary characteristics to improve fusion quality. Unlike prior approaches that overlook the distinct physical properties encoded in frequency spectra, PAD is the first to introduce explicit amplitude-phase decoupling for multi-modal fusion. Specifically, PAD comprises two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features via convolution-guided scaling to enhance geometric consistency; and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive multilayer perceptrons, leveraging SAR’s morphological sensitivity and RGB’s spectral richness. Extensive experiments on WHU-OPT-SAR and DDHR-SK datasets demonstrate state-of-the-art performance. Our work establishes a new paradigm for physics-aware multi-modal fusion in remote sensing. The code will be available at https://github.com/RanFeng2/PAD.
Index Terms:
Land Cover Classification, RGB-SAR multi-modality, multi-modal segmentation, Synthetic Aperture Radar, Remote Sensing
@@ I Introduction
Figure 1: Motivation and SAR-RGB Spatial-Frequency Difference Analysis. (a) SAR and RGB imagery exhibit modality gaps due to distinct sensing mechanisms (backscattering vs. reflectance). (b) Fusion feature space comprises heterogeneous, shared, and complementary components; effective fusion suppresses heterogeneous features while enhancing shared and complementary patterns. (c)-(g) Frequency-domain analysis on WHU-OPT-SAR (see Section III-A): (c) spatial discrepancies exhibit anisotropic distributions; (d) relative amplitude differences concentrate in high-frequency regions; (e) phase discrepancies are minimal and cluster in low-frequency bands; (f)-(g) sampling-rate variations significantly affect spatial discrepancy distributions.
Despite this complementarity, RGB-SAR fusion faces a key challenge: inherent imaging heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB) leads to significant spectral misalignment and hinders effective cross-modal feature integration (see Fig. 1(a)). In SAR images, objects often appear as discrete backscatter responses, lacking fundamental visual cues, such as geometric contours, textures, and consistent shapes, that are readily available in RGB imagery [4]. Compared to natural-image fusion, integrating SAR and RGB data presents pronounced radiometric and geometric disparities, limiting existing frameworks [5, 6, 7].
Deep learning, particularly convolutional neural networks (CNNs), has achieved notable success in LCC by extracting powerful features [8, 9]. To mitigate modality heterogeneity, most approaches use dual-branch architectures to extract multi-scale features independently, then fuse them via element-wise summation, mean aggregation, or channel-wise concatenation before a shared decoder or segmentation head [10]. However, accurate modeling of complementary features is still impeded by SAR-RGB heterogeneity. Inspired by contrastive learning [11, 12, 13], recent works align modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts (see Fig. 1(b)), yet spatial-domain modeling remains inefficient.
@@
We conduct comprehensive spectral analysis (see Fig. 1(c)-(g)) to reveal three critical frequency-domain priors: 1) SAR-RGB spatial-domain differences are anisotropic and susceptible to pixel-level mismatch (see Fig. 1(c)); 2) average phase spectra differ minimally and mainly in low frequencies, indicating phase as shared information (see Fig. 1(e)); and 3) amplitude mismatches exhibit distinct high/low-frequency patterns, most pronounced in the high-frequency band, highlighting the need for frequency-aware modeling (see Fig. 1(d)).
Motivated by these insights, we propose Phase-Amplitude Decoupling (PAD), which disentangles phase components capturing modality-invariant structures from amplitude components encoding modality-specific details. By explicitly aligning shared phase representations, PAD enhances complementary amplitude feature extraction and reduces cross-modal conflicts. PAD comprises two modules: Phase Spectrum Correction (PSC), which aligns and reinforces structural details via convolution-guided scaling; and Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive MLPs.
@@
  • We systematically analyze registration datasets to establish frequency distributions of SAR-RGB common (phase) and unique (amplitude) patterns, offering interpretable physical priors for multi-source fusion.
  • We introduce frequency-domain amplitude-phase decoupling for multi-modal fusion, orthogonally separating phase (modality-shared) and amplitude (modality-specific) features to resolve heterogeneity-induced conflicts.
@@
  • ASF uses frequency-domain MLPs to integrate complementary high-frequency (local details) and low-frequency (global structure) components, leveraging SAR’s morphological sensitivity and RGB’s spectral richness for improved segmentation.
Figure 2: Flowchart of the PAD Framework. Registered RGB and SAR images, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are processed by asymmetric backbones to extract features $x_{\mathrm{RGB}}$ and $x_{\mathrm{SAR}}$ at each stage. These features are passed through $n$ PAD fusion modules, each comprising: (a) the SCF module, (c) the ASF module, and (d) the PSC module. (b) FD and (e) FR denote the spatial-to-frequency and frequency-to-spatial transformations, respectively. Each stage’s fused features are concatenated channel-wise (outer circle $C$) and fed into a shared decoder. Here, $\mathcal{A}$ and $\mathcal{P}$ represent amplitude and phase spectra.
@@ II-A1 Cross-Modal Transfer Learning
Transfer learning [26] adapts knowledge from one modality to another, but the pronounced differences between SAR and RGB data often degrade learned representations, limiting multi-modal applicability.
@@ II-A2 Hierarchical Feature Fusion
Feature fusion has evolved through three stages: a) Basic Fusion Operators: Early methods used concatenation or weighted summation [27], but these are prone to modality dominance and redundancy. b) Attention-Guided Fusion: Spatial-channel attention mechanisms address these issues [28, 29]. For example, MCANet’s cross-modal attention (MCAM) models inter-modality dependencies; MRFS [30] introduces progressive cycle attention for infrared-visible fusion; and CMX [31] employs bidirectional channel-spatial attention. However, these paradigms incur high computational costs. c) Complementarity-Aware Fusion: Recent approaches explicitly disentangle shared and modality-specific features. ASANet [32] uses semantic focusing modules to emphasize complementary cues; CDDFuse [12] applies contrastive learning to separate private and shared representations; and CEN [33] leverages dynamic channel exchange with independent batch normalization. Despite these advances, spatial-domain methods still struggle with geometric misalignment and inefficiency.
@@ II-B Frequency Representation Learning in Vision Tasks
Frequency representation learning has emerged as a powerful tool across various vision tasks due to its capacity for global modeling and structural decomposition. In this subsection, we review three relevant areas:
@@ II-B1 Spectral Decomposition for Structure-Style Separation
Fourier-based decomposition isolates phase (structural semantics) and amplitude (appearance) for domain generalization and style transfer [34, 35]. Early methods used feature statistics for style normalization [36, 37], while recent works employ Fourier transforms for clearer separation, enabling content-preserving augmentation by swapping amplitudes across domains [38, 39].
@@ II-B2 Frequency-Driven Global Modeling
Fourier transforms provide a global receptive field useful for efficient attention and context modeling. Fast Fourier Convolution (FFC) [40] and FFTformer [41] leverage spectral processing to replace or accelerate spatial attention. Spectral token mixing has also been effective in lightweight transformers [42, 43].
@@ II-B3 Frequency-Based Inductive Biases in Dense Prediction
Spectral priors enhance dense prediction tasks. DFF [44] uses frequency-domain dynamic filtering to amplify invariant low-frequency features, improving cross-domain generalization. FreqFusion [45] designs learnable high- or low-pass filters for semantic consistency and boundary refinement. FCENet [14] employs frequency-selective modules for cross-modal spectral correlations in RGB-NIR denoising.
In summary, frequency-based representations offer a flexible, physically interpretable design space for vision tasks; yet, few frameworks target task-specific fusion in remote sensing. We address this gap in the next section.
@@ II-C Frequency-Domain Fusion Methods
Applying frequency representations to multi-modal fusion is recent. Shi et al. [15] introduce Frefusion, using spectral attention to merge modalities in the frequency domain. Frequency spectra naturally decompose signals into phase (modality-invariant structures) and amplitude (modality-specific details), offering a principled fusion framework.
However, existing methods often transplant spatial operators, MLPs [16], convolutions [14], or attention [15], into the spectral domain, modulating magnitude while neglecting the distinct roles of phase and amplitude. Consequently, shared information may be preserved, but complementary cues are suppressed.
In contrast, our PAD framework explicitly decouples phase and amplitude spectra. Through systematic spectral analysis (see Section III-A), we show that SAR and RGB share strong phase consistency while amplitude captures complementary patterns. PAD introduces (1) a Phase Spectrum Correction module for robust structural alignment and (2) an Amplitude Spectrum Fusion module with frequency-adaptive refinement. This decomposition-correction-fusion pipeline distinguishes PAD from previous FFT-based approaches by providing a physically interpretable strategy for multi-modal fusion.
The two most related works, FeINFN [16] and SFINet++ [46], also leverage amplitude and phase in the frequency domain for high-frequency detail enhancement and global context modeling. However, they lack a physically grounded interpretation of amplitude-phase roles and do not explicitly decouple shared versus modality-specific information. In contrast, PAD is the first framework to formulate and exploit amplitude-phase decomposition for principled alignment and complementary integration across modalities.
@@ III Proposed Method
To address the fundamental conflict between modality-invariant geometry and modality-specific appearance in SAR-RGB fusion, we introduce Phase-Amplitude Decoupling (PAD), a frequency-aware multi-modal fusion framework. The overall architecture, shown in Fig. 2, comprises three modules: the Spatial Complementary Fusion (SCF) module, which preserves fine-grained spatial priors; the Phase Spectrum Correction (PSC) module, which aligns and reinforces shared structural details; and the Amplitude Spectrum Fusion (ASF) module, which adaptively integrates complementary high- and low-frequency amplitude cues. Unlike previous methods that apply spectral operations uniformly, PAD explicitly decouples and recombines phase and amplitude with dedicated modules to enable physics-consistent fusion.
@@
2 $I_{\mathrm{RGB}}\leftarrow 0.299R+0.587G+0.114B$
//Retain SAR image as is
@@ 3 $I_{\mathrm{SAR}}\leftarrow I_{\mathrm{SAR}}$
6       Compute $\mathcal{A}_{\mathrm{SAR}}^{(i)}$ and $\mathcal{P}_{\mathrm{SAR}}^{(i)}$ via Eqs. 11, 12 and 13
7       Compute $\mathcal{A}_{\mathrm{RGB}}^{(i)}$ and $\mathcal{P}_{\mathrm{RGB}}^{(i)}$ similarly
@@
      //Spatial difference computation
9       Compute $\Delta I^{(i)}\leftarrow\mathrm{RSD}^{(i)}$ via Eq. 1
       //Amplitude difference computation
10       Compute $\Delta\mathcal{A}^{(i)}\leftarrow\mathrm{RAD}^{(i)}$ via Eq. 2
       //Phase difference computation
11       Compute $\Delta\mathcal{P}^{(i)}\leftarrow\mathrm{APPD}^{(i)}$ via Eq. 3
12
13 end for
14
@@
15 $\mathrm{RSD}\leftarrow\mathtt{Mean}(\{\Delta I^{(i)}\}_{i=1}^{N})$
16 $\mathrm{RAD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{A}^{(i)}\}_{i=1}^{N})$
17 $\mathrm{APPD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{P}^{(i)}\}_{i=1}^{N})$
@@ III-A Frequency-domain Priors for Amplitude-Phase Decoupling
To validate the physical significance of phase-amplitude decoupling and reveal the shared and complementary relations between RGB and SAR, we conducted a systematic spectral analysis on WHU-OPT-SAR dataset, the largest RGB-SAR segmentation benchmark.
Firstly, we visualized the averaged difference spectra (see Fig. 1(c)-(e)) using Algorithm 1 to compare modality differences in spatial, amplitude, and phase domains. This analysis shows that frequency-domain representations capture cross-modal sharing and complementarity more effectively than spatial representations.
@@
$\mathrm{APPD}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\arg\exp j(\mathcal{P}_{\mathrm{SAR}}^{(i)}-\mathcal{P}_{\mathrm{RGB}}^{(i)})\right|,$ (3)
where $I$, $\mathcal{A}$, and $\mathcal{P}$ denote pixel values, amplitude spectra, and phase spectra for SAR and RGB images; $\arg(\cdot)$ computes the complex angle; and $\epsilon$ prevents division by zero. By leveraging the periodicity of complex exponentials, APPD maps phase differences into $[-\pi,\pi)$, correcting wrap-around artifacts (e.g., $2\pi\to 0$).
Figure 1(c)-(e) reveals three priors: 1) Spatial differences (RSD) exhibit anisotropic distributions. 2) Amplitude differences (RAD) concentrate in high-frequency regions. 3) Phase differences (APPD) remain minimal, with slight low-frequency clustering. These frequency-domain metrics demonstrate superior explainability over spatial RSD, confirming that spectral decomposition better captures cross-modal specificity.
Physically, low-frequency phase disparities stem from geometric distortions (e.g., side-looking SAR vs. orthographic RGB), while high-frequency variations arise from SAR speckle noise. Amplitude disparities reflect modality-specific sensitivity to object scales: low frequencies encode large structures (farmlands, forests) and high frequencies capture fine details (roads).
These analyses yield two key insights that drive our architecture design: 1) Phase Sharing: Consistently low APPD values indicate strong phase consistency. The PSC module thus performs pixel-level alignment while preserving this shared information. 2) Amplitude Complementarity: RAD’s frequency-dependent divergence motivates ASF’s dynamic frequency-radius adjustment to handle amplitude variations.
Secondly, we conducted sensitivity experiments on sampling rates (see Fig. 1(f)-(g)) to evaluate the stability of our spectral metrics. By downsampling input images at various ratios and computing the mean and variance of RSD, RAD, and APPD across the dataset, we observed that RAD and APPD remain nearly constant, whereas RSD exhibits larger fluctuations. This invariance of RAD and APPD to sampling artifacts confirms that frequency-domain complementarity arises from intrinsic modality characteristics rather than from individual changes.
TABLE I: Statistical Analysis of the Average Periodic Phase Difference Spectrum.
Metric Low Frequency High Frequency Whole Image
@@
Figure 3: Statistical Analysis of APPD. (a) APPD spectrum. (b)-(d) Frequency histograms with Gaussian fits for low-frequency (LF; radius = 0.5 $\times$ half-diagonal), high-frequency (HF), and all spectra (ALL). (e)-(f) Mean-variance trajectories across sampling rates. All spectra exhibit near-Gaussian invariance, with LF and ALL showing slight deviations due to structural noise in the low-frequency range, indicating that phase, particularly HF, can serve as a robust shared feature.
Thirdly, to assess the statistical consistency of cross-modal phase spectra, we performed normality tests (Shapiro-Wilk) and calculated skewness and kurtosis for APPD values at each frequency band (see Table I). Frequency histograms and mean-variance trajectories (see Fig. 3) demonstrate that APPD distributions closely follow a Gaussian profile, especially in high-frequency regions. Low-frequency components show minor deviations, likely due to structural noise, but overall the near-Gaussian behavior validates phase spectra as a reliable shared feature. In a word, phase is so similar between RGB and SAR that it is almost unaffected by individual differences.
Figure 4: Visualization of Spatial and Frequency Spectrum Differences under Missing RGB Regions. (a)-(b) RGB images with missing regions; (c) variation in RSD, RAD, and APPD, showing RSD’s greater sensitivity to missing data; (d)-(f) difference spectra including missing regions; (g)-(i) difference spectra excluding missing regions. These results demonstrate that frequency-domain metrics (RAD, APPD) are more robust to missing data than spatial-domain RSD.
Figure 5: Visualization of Spatial and Frequency Spectrum Differences under Cloud Obscuration. (a)-(c) Example observations with cloud masks derived from RGB, SAR, and NIR channels in WHU-OPT-SAR [47]. (d) Comparison of RSD, RAD, and APPD stability between cloud-covered and cloud-free areas; arrows denote the value of decline.
Finally, we evaluated phase invariance under two challenging scenarios: 1) partial RGB data removal (see Fig. 4) and 2) cloud-obscured observations (see Fig. 5). In the former, RSD increased dramatically when RGB pixels were missing (compare Fig. 4(d) and (g)), while APPD remained stable (compare Fig. 4(f) and (i)). In the latter, APPD variance ($\Delta\approx 10^{-0.037}\%$) was negligible compared to that of RSD ($\Delta\approx 10^{5.81}\%$) and RAD ($\Delta\approx 10^{0.58}\%$) as shown in Fig. 5(d). These results confirm that phase information, particularly at high frequencies, retains robust consistency across modalities, underscoring its role as a shared feature to mitigate inter-modal conflicts and enhance complementary fusion.

### v3

@@ Abstract
The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and underexploited spectral complementarity. Existing approaches often fail to decouple shared structural features from modality-complementary radiometric attributes, resulting in feature conflicts and information loss. To address this, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-complementary) components in the Fourier domain. This design reinforces shared structures while preserving complementary characteristics, thereby enhancing fusion quality. Unlike previous methods that overlook the distinct physical properties encoded in frequency spectra, PAD explicitly introduces amplitude-phase decoupling for multi-modal fusion. Specifically, PAD comprises two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features via convolution-guided scaling to improve geometric consistency; and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high- and low-frequency patterns using frequency-adaptive multilayer perceptrons, effectively exploiting SAR’s morphological sensitivity and RGB’s spectral richness. Extensive experiments on WHU-OPT-SAR and DDHR-SK demonstrate state-of-the-art performance. This work establishes a new paradigm for physics-aware multi-modal fusion in remote sensing. The code will be available at https://github.com/RanFeng2/PAD.
@@ I Introduction
Figure 1: Motivation and SAR-RGB Spatial-Frequency Difference Analysis. (a) SAR and RGB imagery exhibit modality gaps due to distinct sensing mechanisms (backscattering vs. reflectance). (b) The fusion feature space contains heterogeneous, shared, and complementary components; effective fusion suppresses heterogeneous cues while enhancing shared and complementary patterns. (c-g) Frequency-domain analysis on WHU-OPT-SAR (see §III-A): (c) spatial discrepancies show anisotropic distributions; (d) relative amplitude differences concentrate in high-frequency regions; (e) phase discrepancies are minimal and primarily cluster in low-frequency bands; (f-g) varying the sampling rate significantly affects spatial discrepancy distributions.
Despite this complementarity, RGB-SAR fusion faces a key challenge: the inherent heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB) causes spectral misalignment and hinders effective cross-modal feature integration (see Fig. 1(a)). In SAR images, objects often appear as discrete backscatter responses, lacking fundamental visual cues, such as geometric contours, textures, and consistent shapes, that are readily available in RGB imagery [4]. Compared with natural image fusion, integrating SAR and RGB data presents pronounced radiometric and geometric disparities, which limits existing frameworks [5, 6, 7].
Deep learning, particularly convolutional neural networks (CNNs), has achieved remarkable success in LCC by extracting powerful features [8, 9, 10, 11]. To mitigate modality heterogeneity, most approaches employ dual-branch architectures to extract multi-scale features independently and then fuse them via element-wise summation, mean aggregation, or channel-wise concatenation before a shared decoder or segmentation head [12]. However, accurately modeling complementary features remains impeded by SAR-RGB heterogeneity. Inspired by contrastive learning [13, 14], recent works align modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts (see Fig. 1(b)), yet spatial-domain modeling remains inefficient.
@@
We conduct a comprehensive spectral analysis (see Fig. 1(c)-(g)) to reveal three critical frequency-domain priors: 1) SAR-RGB spatial-domain differences are anisotropic and susceptible to pixel-level mismatch (see Fig. 1(c)); 2) average phase spectra differ minimally and primarily in low frequencies, indicating phase as shared information (see Fig. 1(e)); and 3) amplitude mismatches exhibit distinct high- and low-frequency patterns, most pronounced in the high-frequency band, highlighting the need for frequency-aware modeling (see Fig. 1(d)).
Motivated by these insights, we propose Phase-Amplitude Decoupling (PAD), which decouples phase components capturing modality-shared structures from amplitude components encoding modality-complementary details. By explicitly aligning shared phase representations, PAD enhances complementary amplitude feature extraction and reduces cross-modal conflicts. PAD comprises two modules: Phase Spectrum Correction (PSC), which aligns and reinforces structural details via convolution-guided scaling; and Amplitude Spectrum Fusion (ASF), which dynamically integrates high- and low-frequency patterns using frequency-adaptive multilayer perceptrons (MLPs).
@@
  • We systematically analyze registered datasets to establish the frequency distributions of SAR-RGB shared (phase) and complementary (amplitude) patterns, providing interpretable physical priors for multi-source fusion.
  • We introduce frequency-domain amplitude-phase decoupling for multi-modal fusion, orthogonally separating phase (modality-shared) and amplitude (modality-complementary) features to resolve heterogeneity-induced conflicts.
@@
  • ASF leverages frequency-adaptive MLPs to integrate complementary high-frequency (local details) and low-frequency (global structure) components, exploiting SAR’s morphological sensitivity and RGB’s spectral richness for improved segmentation.
Figure 2: Flowchart of the PAD Framework. Registered RGB and SAR inputs, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are processed by asymmetric backbones to extract features $x_{\mathrm{RGB}}$ and $x_{\mathrm{SAR}}$ at each stage. These features pass through $n$ PAD fusion modules, each comprising: (a) SCF; (b) FD (spatial-to-frequency transform); (c) ASF; (d) PSC; and (e) FR (frequency-to-spatial transform). At each stage, the fused features are concatenated channel-wise (outer circle, $C$) and fed into a shared decoder. Here, $\mathcal{A}$ and $\mathcal{P}$ denote the amplitude and phase spectra, respectively.
@@ II-A1 Cross-Modal Transfer Learning
Transfer learning [27] adapts knowledge from one modality to another, but the pronounced differences between SAR and RGB often degrade learned representations, limiting their multi-modal applicability.
@@ II-A2 Hierarchical Feature Fusion
Feature fusion has evolved through three stages: a) Basic Fusion Operators: Early methods employed concatenation or weighted summation [28], but these are prone to modality dominance and redundancy. b) Attention-Guided Fusion: Spatial-channel attention mechanisms alleviate these issues [29, 30]. For example, MCANet’s cross-modal attention (MCAM) models inter-modality dependencies; MRFS [31] introduces progressive cycle attention for infrared-visible fusion; and CMX [32] applies bidirectional channel-spatial attention. However, these paradigms incur high computational costs. c) Complementarity-Aware Fusion: Recent approaches explicitly decouple shared and complementary features. ASANet [33] uses semantic focusing modules to emphasize complementary cues; CDDFuse [13] applies contrastive learning to separate private and shared representations; and CEN [34] leverages dynamic channel exchange with independent batch normalization. Despite these advances, spatial-domain methods still struggle with geometric misalignment and inefficiency.
@@ II-B Frequency Representation Learning in Vision Tasks
Frequency representation learning has emerged as a powerful tool across various vision tasks due to its capacity for global modeling and structural decomposition. We briefly review three relevant areas:
@@ II-B1 Spectral Decomposition for Structure-Style Separation
Fourier-based decomposition isolates phase (structural semantics) and amplitude (appearance) for domain generalization and style transfer [35]. Early methods relied on feature statistics for style normalization [36, 37], while recent works apply Fourier transforms for clearer separation, enabling content-preserving augmentation by swapping amplitudes across domains [38, 39].
@@ II-B2 Frequency-Driven Global Modeling
Fourier transforms provide a global receptive field useful for efficient attention and context modeling. Fast Fourier Convolution (FFC) [40] and FFTformer [41] leverage spectral processing to replace or accelerate spatial attention. Spectral token mixing has also proven effective in lightweight transformers [42, 43].
@@ II-B3 Frequency-Based Inductive Biases in Dense Prediction
Spectral priors benefit dense prediction tasks [44, 45, 46, 47]. DFF [48] uses frequency-domain dynamic filtering to amplify invariant low-frequency features, improving cross-domain generalization. FreqFusion [49] designs learnable high- and low-pass filters for semantic consistency and boundary refinement. FCENet [15] employs frequency-selective modules to model cross-modal spectral correlations in RGB-NIR denoising.
In summary, frequency-based representations offer a flexible and physically interpretable design space for vision tasks. Yet, few frameworks address task-specific fusion in remote sensing. We target this gap in the next section.
@@ II-C Frequency-Domain Fusion Methods
The application of frequency representations to multi-modal fusion is relatively recent. Shi et al. [16] introduce Frefusion, which applies spectral attention to merge modalities in the frequency domain. Frequency spectra naturally decompose signals into phase (modality-shared structures) and amplitude (modality-complementary details), offering a principled basis for fusion.
However, existing methods often transplant spatial operators, such as MLPs [17], convolutions [15], or attention [16], into the spectral domain, modulating magnitude while neglecting the distinct roles of phase and amplitude. As a result, shared information may be preserved, but complementary cues are frequently suppressed.
In contrast, our PAD framework explicitly decouples phase and amplitude spectra. Through systematic spectral analysis (see §III-A), we show that SAR and RGB exhibit strong phase consistency, while amplitude captures complementary patterns. PAD introduces 1) a Phase Spectrum Correction (PSC) module, which performs robust structural alignment, and 2) an Amplitude Spectrum Fusion (ASF) module, which conducts frequency-adaptive refinement. This decomposition-correction-fusion pipeline distinguishes PAD from previous FFT-based approaches by providing a physically interpretable strategy for multi-modal fusion.
The two most related works, FeINFN [17] and SFINet++ [50], also leverage amplitude and phase in the frequency domain for high-frequency detail enhancement and global context modeling. However, they lack a physically grounded interpretation of amplitude-phase roles and do not explicitly decouple shared versus complementary information. In contrast, PAD is the first framework to formulate and exploit amplitude-phase decomposition for principled alignment and complementary integration across modalities.
@@ III Proposed Method
To address the fundamental conflict between modality-shared geometry and modality-complementary appearance in SAR-RGB fusion, we introduce Phase-Amplitude Decoupling (PAD), a frequency-aware multi-modal fusion framework. The overall architecture in Fig. 2 comprises three modules: the Spatial Complementary Fusion (SCF) module, which preserves fine-grained spatial priors; the Phase Spectrum Correction (PSC) module, which aligns and reinforces shared structural details; and the Amplitude Spectrum Fusion (ASF) module, which adaptively integrates complementary high- and low-frequency amplitude cues. Unlike previous methods that apply spectral operations uniformly, PAD explicitly decouples and then recombines phase and amplitude with dedicated modules for physics-consistent fusion.
@@
$I_{\mathrm{RGB}}\leftarrow 0.299R+0.587G+0.114B$ //Retain SAR image as is
@@ 3 $I_{\mathrm{SAR}}\leftarrow I_{\mathrm{SAR}}$
5  Compute $\mathcal{A}_{\mathrm{SAR}}^{(i)}$ and $\mathcal{P}_{\mathrm{SAR}}^{(i)}$ via 2D FFT and FFTshift; Compute $\mathcal{A}_{\mathrm{RGB}}^{(i)}$ and $\mathcal{P}_{\mathrm{RGB}}^{(i)}$ similarly
 //Spatial difference computation
  Compute $\Delta I^{(i)}\leftarrow\mathrm{RSD}^{(i)}$ via Eq. 1 //Amplitude difference computation
  Compute $\Delta\mathcal{A}^{(i)}\leftarrow\mathrm{RAD}^{(i)}$ via Eq. 2 //Phase difference computation
6  Compute $\Delta\mathcal{P}^{(i)}\leftarrow\mathrm{APPD}^{(i)}$ via Eq. 3
7 end for
@@
$\mathrm{RSD}\leftarrow\mathtt{Mean}(\{\Delta I^{(i)}\}_{i=1}^{N})$ $\mathrm{RAD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{A}^{(i)}\}_{i=1}^{N})$ $\mathrm{APPD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{P}^{(i)}\}_{i=1}^{N})$
@@ III-A Frequency-Domain Priors for Amplitude-Phase Decoupling
To validate the physical significance of amplitude-phase decoupling and reveal the shared and complementary relations between RGB and SAR, we conduct a systematic spectral analysis on WHU-OPT-SAR, the largest RGB-SAR segmentation benchmark.
First, we visualize averaged difference spectra (see Fig. 1(c)-(e)) using Algorithm 1 to compare modality differences in the spatial, amplitude, and phase domains. This analysis shows that frequency-domain representations capture cross-modal sharing and complementarity more effectively than spatial representations.
@@
$\mathrm{APPD}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\arg\exp\left(\mathrm{i}(\mathcal{P}_{\mathrm{SAR}}^{(i)}-\mathcal{P}_{\mathrm{RGB}}^{(i)})\right)\right|,$ (3)
where $I$ denotes pixel intensities; $\mathcal{A}$ and $\mathcal{P}$ are amplitude and phase spectra, respectively; $\arg(\cdot)$ returns the complex angle; $\mathrm{i}=\sqrt{-1}$; and $\epsilon>0$ prevents division by zero. By leveraging the periodicity of complex exponentials, Eq. 3 maps phase differences to $\pi,\pi)$, correcting wrap-around artifacts (e.g., $2\pi\to 0$).
Figure 1(c)-(e) reveals three priors: 1) spatial differences (RSD) exhibit anisotropic distributions; 2) amplitude differences (RAD) concentrate in high-frequency regions; and 3) phase differences (APPD) remain minimal, with slight low-frequency clustering. These frequency-domain metrics provide better explainability than spatial RSD, confirming that spectral decomposition more faithfully captures cross-modal characteristics.
Physically, low-frequency phase disparities stem from geometric distortions (e.g., side-looking SAR vs. orthographic RGB), while high-frequency variations arise from SAR speckle. Amplitude disparities reflect modality-complementary sensitivity to object scales: low frequencies encode large structures (farmlands, forests), whereas high frequencies capture fine details (roads).
These analyses yield two key insights that guide our design: Phase sharing. Consistently low APPD indicates strong phase consistency; PSC therefore performs pixel-level alignment while preserving this shared information. Amplitude complementarity. The frequency-dependent divergence of RAD motivates ASF’s dynamic frequency-radius adjustment to handle amplitude variations.
Second, we conduct sensitivity experiments on sampling rates (see Fig. 1(f)-(g)). By downsampling images at various ratios and computing the mean/variance of RSD, RAD, and APPD across the dataset, we observe that RAD and APPD remain nearly constant, whereas RSD fluctuates more. This invariance indicates that frequency-domain complementarity arises from intrinsic modality characteristics rather than sampling artifacts.
TABLE I: Statistical Analysis of the Average Periodic Phase Difference (APPD) Spectrum.
Metric
Low frequency
High frequency
ALL
@@
Figure 3: Statistical Analysis of APPD. (a) APPD spectrum. (b-d) Frequency histograms with Gaussian fits for low frequency (LF; radius = $0.5\times$ half-diagonal), high frequency (HF), and all spectra (ALL). (e-f) Mean-variance trajectories across sampling rates. Across sampling rates, all spectra exhibit near-Gaussian behavior, while slight deviations in LF and ALL are attributable to structural noise at low frequencies, indicating that the phase, particularly HF, serves as a robust shared feature.
Third, we assess statistical consistency of cross-modal phase spectra via Shapiro-Wilk tests and by computing skewness and kurtosis of APPD within each frequency band (see Table I). Frequency histograms and mean-variance trajectories (see Fig. 3) show that APPD distributions closely follow a Gaussian profile, especially in high-frequency regions; small low-frequency deviations likely result from structural noise. In summary, RGB and SAR phases are highly similar and largely unaffected by instance-level variations.
Figure 4: Visualization of Spatial-Frequency Spectrum Differences under Missing RGB Regions. (a-b) Two RGB images have missing regions in the original dataset. (c) Variation in RSD, RAD, and APPD, indicating RSD’s greater sensitivity to missing data. (d-f) Difference spectra computed including the missing regions. (g-i) Difference spectra computed excluding the missing regions. Overall, frequency-domain metrics (RAD and APPD) are more robust to missing data than the spatial-domain RSD.
Figure 5: Visualization of Spatial-Frequency Spectrum Differences under Cloud Obscuration. (a-c) Example observations with cloud masks derived from the RGB, SAR, and NIR channels in WHU-OPT-SAR [51], respectively. (d) Comparison of RSD, RAD, and APPD between cloud-covered and cloud-free areas; arrows indicate the value of the decrease.
Finally, we evaluate phase invariance under two challenging scenarios: 1) partial RGB removal (see Fig. 4), and 2) cloud occlusion (see Fig. 5). In the former, RSD increases markedly when RGB pixels are missing (see Fig. 4(d) and (g)), whereas APPD remains stable (see Fig. 4(f) and (i)). In the latter, the variance of APPD ($\Delta\approx 10^{-0.037}$) is negligible compared with that of RSD ($\Delta\approx 10^{5.81}$) and RAD ($\Delta\approx 10^{0.58}$) in Fig. 5(d). These results confirm that phase, particularly at high frequencies, retains robust cross-modal consistency, underscoring its role as a shared feature to mitigate inter-modal conflicts and enhance modality-complementary fusion.